# A Foundational Theory for Decentralized Sensory Learning

### v1

@@ Abstract
In both neuroscience and artificial intelligence, popular functional frameworks and neural network formulations operate by making use of extrinsic error measurements and global learning algorithms. Through a set of conjectures based on evolutionary insights on the origin of cellular adaptive mechanisms, we reinterpret the core meaning of sensory signals to allow the brain to be interpreted as a negative feedback control system, and show how this could lead to local learning algorithms without the need for global error correction metrics. Thereby, a sufficiently good minima in sensory activity can be the complete reward signal of the network, as well as being both necessary and sufficient for biological learning to arise. We show that this method of learning was likely already present in the earliest unicellular life forms on earth. We show evidence that the same principle holds and scales to multicellular organisms where it in addition can lead to division of labour between cells. Available evidence shows that the evolution of the nervous system likely was an adaptation to more effectively communicate intercellular signals to support such division of labour. We therefore propose that the same learning principle that evolved already in the earliest unicellular life forms, i.e. negative feedback control of externally and internally generated sensor signals, has simply been scaled up to become a fundament of the learning we see in biological brains today. We illustrate diverse biological settings, from the earliest unicellular organisms to humans, where this operational principle appears to be a plausible interpretation of the meaning of sensor signals in biology, and how this relates to current neuroscientific theories and findings.
@@ 1.3 Multicellular evolution
Therefore, in an emergent strategy, early multicellular organisms built highly complex networks of chemical signaling (Figure 3A,B), and many of these diverse chemical signals, i.e. neuropeptides and other low-molecular weight compounds, are still used as neurotransmitters by the hypothalamus in animals today Arendt [2008], Jékely et al. [2015]. This is an extension of the same strategy already present within individual cells as discussed above, i.e. where a chemical process triggered by a sensory signal results in an internal chemical problem signal that impacts the molecular cascade reactions within the same cell. In multicellular organisms, those chemical problem signals can in addition be passed externally, between the cells, representing the request of cell A to solve a problem that it does not have the capacity to solve by itself (see Figure 3). This could naturally happen also in colonies of homogenous cell types. The advantage of multiple differentiated cells is that the set of responses, or problem solutions, that the organism is capable of can be made more diversified and scaled to enact larger changes in the surroundings. Incorporating these points, we propose the following conjecture:
@@
In many circumstances, this gives rise to secondary effects: Two, or more, competing problems can through their transmission give rise to an externally observable action that resolves one of the problems and simultaneously a prolonged activity within the organism to retain the others, which in this period may very well grow larger. Such a phenomena can give rise to situations where the activity increases in one part of the organism as one problem is retained - or even created - in the process of solving another, more acute problem.
@@ 1.4 Neurons and Networks
The evolution of neural cells, and possibly the divergence between plant-like and animal-like organisms, is also closely linked to the evolution of cells with contractile properties Jékely et al. [2015], Arendt [2021], which enabled the speed-up of ingestion due to passage of sea water into the digestive canal, but also movement of the organism as a whole towards better nutrient sources, as exemplified by the hydra Szymanski and Yuste [2019], Wang et al. [2023]. Hence, the multicellular organisms were now provided with another ‘tool’ of actuation, even though contractile cells or muscle cells just represent a further differentiation of the mechanoactuation capability provided by the cytoskeleton present already in the generic cell (Figure 1), with which the organism could solve its problems of nutrient supply.
@@ 1.5 Sensory Prioritization and Negative Feedback in organisms and animals
Indeed, the sensory afferent systems of the skin, ligaments and muscles are arranged in terms of falling activation thresholds, where the sensors of the highest mechanical threshold are chemosensors (group IV muscle afferents Amann et al. [2015], C fiber afferents in the skin Mense [1996] signifying that the mechanical load, for example in the skin, has become so large as to tear the cells of the local tissue apart. In fact, even the skin sensors of the lowest mechanical threshold are in essence shear force sensors Hayward et al. [2014], Jörntell et al. [2014], Rongala et al. [2024], implying that every sensor type in the skin to some extent forebodes tissue tear but to different hierarchical degrees, in a way that can support a sensory prioritization schema by the nervous system.
@@
What we suggest and further exemplify below is that the acuteness of different sensory signals can be comparable across modalities if they reflect an equally critical problem. When one problem or threat is greater than another, that is the dominating factor which gains control over the system. In this case, sensors that are typically perceived as being used to make predictions can in some contexts suddenly become of a more acute nature and other sensory systems which in the default situation are more critical threat signals can then almost be seen as becoming supportive. Importantly, the behaviour of the sensor hasn’t changed, only the importance and criticality of the problem it perceives.
@@
It should be clear here that when describing a problem signal, such a signal can be transmitted through the body in many forms. It could be in a chemical signal form like glutamate, GABA, or some monoamine or peptide, through the neural system or through the bloodstream. In the neural system, it could be in an electrical form, binary as with an action potential or analog as a hyperpolarizing or depolarizing membrane voltage deviation from its resting value, i.e. set point. It could even be in the form of a chemical imbalance throughout the body, like an overabundance or shortage of calcium. This also points at how such problems are resolved, as a hyperpolarizing signal could be balanced with a depolarizing signal to reduce its effect in some parts of the body, whereas in other parts it could end up resulting in a muscle contraction, or an overabundance of calcium could be solved by opening channels which moves such calcium into intracellular buffer systems, whereas a calcium-deprived state could be resolved by releasing calcium from the same buffer system.
@@ 1.6 Relationship to classical feedback control
Figure 4: Principle of operation of a PID system in negative feedback control. A set point of the loop is compared to the actual current sensor signal. A deviation constitutes a ‘problem’ that needs to be fixed. The problem signal is fed through a converter that weighs in proportional (P) gain of the signal, an integration (I) of the signal and a derivation (D) of the signal to create an actuation control signal. The resulting actuation effect, i.e. depending on the actuation mechanisms that the cells have affordances for and the resulting impact on the environment, is sensed and fed back to the summator. In biology, the PID effects could be other types of transformation effects, or representations of the problem signal, that the cells and the cellular network is capable of, as described in the first section.
@@
Hence, in this article we propose neuronal networks to have only sensory inputs which are increasing in relation to their measured deviation from a set point; i.e. an error for that parameter. These networks would exclude purely informative sensory inputs altogether unless they can be ascribed to a set point and error vector before being introduced to the network, and as observed even tactile sensors detecting shear forces, or retinal cones detecting light changes can be seen in this perspective. For example, the 50,000-100,000 tactile skin sensors in humans would then primarily be seen as ultrasensitive skin shear force sensors Hayward et al. [2014], Jörntell et al. [2014], Rongala et al. [2024], sensing risk of cell damage. The fact that the brain is able to use these lower threshold sensors for more intelligent behavior naturally doesn’t contradict that cell damage risk sensing is the original reason they are there.
@@ 1.7 Relation to Learning Frameworks and Plasticity
It is our hope that the framework we provide here can help explain leading neuroscience theories and experimental observations on biological learning, for example why Hebbian learning and spike timing dependent plasticity are inferrable in experiments, while networks over time still arrive at highly stable homeostatic states. It should be noted that observations that unicellular organisms learn dates back more than 100 years (reviewed by Gershman et al 2021, Gershman et al. [2021]) and the idea that learning in the brain can be seen as a result of the need to keep the individual cells in homeostatic balance was proposed before Levin [2019]. We build on and extend those ideas towards more specific mechanistic explanations through formalizing what is being sensed, how it is being sensed and prioritized, and the passing of locally unresolved problems as internally created sensor signals to more suitable subspaces.
@@
In Artificial Intelligence, input is typically in the form of structured data, to be transformed with the intent of extracting more valuable data. By supplying large amounts of such input to these networks and utilizing training to translate that data into valuable information for a target task, the field of deep learning builds intelligent systems. We want to applaud the impressive results in that field, but here we aimed to show through the presented conjectures that biological organisms very likely do not share the same process and most probably cannot intrinsically value raw data in the form that deep learning makes use of it. Instead, biological organisms sense only solvable, value-estimable problems that need to be surmounted to ensure survival, and that limitation is what allows those organisms to learn in the first place.

### v2

@@ Abstract
In both neuroscience and artificial intelligence, popular functional frameworks and neural network formulations operate by making use of extrinsic error measurements and global learning algorithms. Through a set of conjectures based on evolutionary insights on the origin of cellular adaptive mechanisms, we reinterpret the core meaning of sensory signals to allow the brain to be interpreted as a negative feedback control system, and show how this could lead to local learning algorithms without the need for global error correction metrics. Thereby, a sufficiently good minima in sensory activity can be the complete reward signal of the network, as well as being both necessary and sufficient for biological learning to arise. We show that this method of learning was likely already present in the earliest unicellular life forms on earth. We show evidence that the same principle holds and scales to multicellular organisms where it in addition can lead to division of labour between cells. Available evidence shows that the evolution of the nervous system likely was an adaptation to more effectively communicate intercellular signals to support such division of labour. We therefore propose that the same learning principle that evolved already in the earliest unicellular life forms, i.e. negative feedback control of externally and internally generated sensor signals, has simply been scaled up to become a fundament of the learning we see in biological brains today. We illustrate diverse biological settings, from the earliest unicellular organisms to humans, where this operational principle appears to be a plausible interpretation of the meaning of sensor signals in biology, how this relates to current neuroscientific theories and findings, and how it can be applied to solve body control.
@@ 1.3 Multicellular evolution
Therefore, in an emergent strategy, early multicellular organisms built highly complex networks of chemical signaling (Figure 3A,B), and many of these diverse chemical signals, i.e. neuropeptides and other low-molecular weight compounds, are still used as neurotransmitters by the hypothalamus in animals today Arendt [2008], Jékely et al. [2015]. This is an extension of the same strategy already present within individual cells as discussed above, i.e. where a chemical process triggered by a sensory signal results in an internal chemical problem signal that impacts the molecular cascade reactions within the same cell. In multicellular organisms, those chemical problem signals can in addition be passed externally, between the cells, representing the request of a cell to solve a problem that it does not have the capacity to solve by itself (see Figure 3). This could naturally happen also in colonies of homogenous cell types. The advantage of multiple differentiated cells is that the set of responses, or problem solutions, that the organism is capable of can be made more diversified and scaled to enact larger changes in the surroundings. Incorporating these points, we propose the following conjecture:
@@
In many circumstances, this gives rise to secondary effects: Two, or more, competing problems can through their transmission give rise to an externally observable action that resolves one of the problems and simultaneously a prolonged activity within the organism to retain the others, which in this period may very well grow larger. Such phenomena can give rise to situations where the activity increases in one part of the organism as one problem is retained - or even created - in the process of solving another, more acute problem.
@@ 1.4 Neurons and Networks
The evolution of neural cells, and possibly the divergence between plant-like and animal-like organisms, is also closely linked to the evolution of cells with contractile properties Jékely et al. [2015], Arendt [2021]. As organisms increased in size, more efficient energy intake became necessary to assist the specialized cells within the larger organism, and subsequently more effective means of propulsion as nutrients were depleted faster. The evolution of cells with contractile properties enabled both the speed-up of ingestion due to passage of sea water into the digestive canal, and movement of the organism as a whole towards better nutrient sources, as exemplified by the hydra Szymanski and Yuste [2019], Wang et al. [2023]. Hence, the multicellular organisms were now provided with another ‘tool’ of actuation, even though contractile cells or muscle cells just represent a further differentiation of the mechanoactuation capability provided by the cytoskeleton present already in the generic cell (Figure 1), with which the organism could solve its problems of nutrient supply.
@@ 1.5 Sensory Prioritization and Negative Feedback in organisms and animals
Indeed, the sensory afferent systems of the skin, ligaments and muscles are arranged in terms of falling activation thresholds, where the sensors of the highest mechanical threshold are chemosensors (group IV muscle afferents Amann et al. [2015], C fiber afferents in the skin Mense [1996]) signifying that the mechanical load, for example in the skin, has become so large as to tear the cells of the local tissue apart. In fact, even the skin sensors of the lowest mechanical threshold are in essence shear force sensors Hayward et al. [2014], Jörntell et al. [2014], Rongala et al. [2024], implying that every sensor type in the skin to some extent forebodes tissue tear but to different hierarchical degrees, in a way that can support a sensory prioritization schema by the nervous system.
@@
What we suggest and further exemplify below is that the acuteness of different sensory signals can be comparable across modalities if they reflect an equally critical problem. When one problem or threat is greater than another, that is the dominating factor which gains control over the system. In this case, sensors that are typically perceived as being used to make predictions can in some contexts suddenly become of a more acute nature and other sensory systems which in the default situation are more critical threat signals can then almost be seen as becoming supportive. Importantly, the behaviour of the sensor hasn’t changed, only the importance and criticality of the problem it monitors.
@@
It should be clear here that a problem signal can be transmitted through the body in many forms. It could be in a chemical signal form like glutamate, GABA, or some monoamine or peptide, through the neural system or through the bloodstream. In the neural system, it could be in an electrical form, binary as with an action potential or analog as a hyperpolarizing or depolarizing membrane voltage deviation from its resting value, i.e. set point. It could even be in the form of a chemical imbalance throughout the body, like an overabundance or shortage of calcium. This also points at how such problems are resolved, as a hyperpolarizing signal could be balanced with a depolarizing signal to reduce its effect in some parts of the body, whereas in other parts it could end up resulting in a muscle contraction, or an overabundance of calcium could be solved by opening channels which moves such calcium into intracellular buffer systems, whereas a calcium-deprived state could be resolved by releasing calcium from the same buffer system.
@@ 1.6 Relationship to classical feedback control
Figure 4: Principle of operation of a PID system in negative feedback control. A set point of the loop is compared to the actual current sensor signal. A deviation constitutes a ‘problem’ that needs to be fixed. The problem signal is fed through a converter that weighs in proportional (P) gain of the signal, an integration (I) of the signal and a derivation (D) of the signal to create an actuation control signal. The resulting actuation effect, i.e. depending on the actuation mechanisms that the cells have affordances for and the resulting impact on the environment, is sensed and fed back to the summator. In biology, the PID effects could be caused by various transformations, or representations of the problem signal, moreover in nested structures, that the cells and the networks of cells are capable of, as described in the Section 1.2.
@@
Hence, in this article we propose neuronal networks to have only sensory inputs which are increasing in relation to their measured deviation from a set point; i.e. an error for that parameter. These networks would exclude purely informative sensory inputs altogether unless they can be ascribed to a set point and error vector before being introduced to the network, and as observed even tactile sensors detecting shear forces, or retinal cones detecting light changes can be seen in this perspective. For example, the 50,000-100,000 tactile skin sensors in humans would then primarily be seen as ultrasensitive skin shear force sensors Hayward et al. [2014], Jörntell et al. [2014], Rongala et al. [2024], sensing risk of cell damage, or simply cell strain. The fact that the brain is able to use these lower threshold sensors for more intelligent behavior naturally doesn’t contradict that sensing the risk of cell damage is the original reason they are there.
@@
1.7 Cell learning applied to problem solving
We next illustrate how the cell learning principles described above can be applied to solve problems in an example control task. It should be noted that this is a bare-bones model to illustrate how a neural or even a non-neural cell system could be constructed from our conjectures. While optimizations are surely possible they would bring an increase in the model complexity and are not needed for this illustration. For example inhibitory neurons, or cells whose output inhibits the activity of other cells, which could increase the behavioral repertoire of the system Spanne and Jörntell [2015], are not included. The system implementations below share similarities with the neuronal circuitry of the spinal cord Jankowska [1992], Kohler et al. [2022]. But note that the systems illustrated below could in principle alternatively be implemented through adaptations in nested molecular cascade reactions (see Section 1.2) in an individual, non-neural cell.
For this illustration, we chose the classical control problem of balancing an inverted pendulum (Figure 5A), in which a pole is balanced on a cart that moves along a rail. The controller can only act by applying forces to the cart with the goal of maintaining the pole in an upright balance. The simulation of the inverted pendulum was conducted in a third party physics simulator, MuJoCo Todorov et al. [2012]. In the present implementation, there were two target set points for the control, the angle of the pole (in one plane only) and the position of the cart along the rail. Success was when these set points were maintained within a range close to zero for a sustained duration of several tens of seconds. To enable control, there were two main categories of sensors, those related to the angle of the pole and those related to the position of the cart. For each category of sensors, there was one sensor signal for each of the two directions and then in addition one sensor signal for the respective derivatives of those signals, yielding a total of eight sensors (Figure 5B). The sensor inputs, i.e. the outputs of the sensor cells, which in a neural system would correspond to its primary sensory afferents, were provided to two actuator cells ($M_{0}$, $M_{1}$; which could be said to correspond to alpha-motorneurons in the spinal cord). The actuator cells pushed the cart with a force to either the right or to the left, which depended on their relative activation levels (Figure 5B).
1.7.1 Implementation of the decentralized sensory learning process
For each cell that received synaptic inputs, which in the first configuration (Figure 5B) only applied to the two output cells, the following set of equations describes how the synaptic inputs were adapted as a function of their activation and the output of the cell (’cells’ will henceforth be referred to as ’neurons’):
$A=\sum_{i=1}^{n}a_{i}.w_{i}$ (1)
$\Delta w_{i}=\begin{cases}\eta,&\text{if }a_{i}>2A\\ -\eta,&\text{if }A>1\\ 0,&\text{if }a_{i}>2A~\&~A>1\end{cases}$ (2)
In addition to weight changes, the network was also adapting its structure over time. If a neuron had activity (A) greater than 1 at the same time as the input of an individual synapse ($a_{i}$) to that neuron was greater than 1, then that synaptic connection was disconnected and reinserted randomly to eligible recipient neurons (see Appendix 1). This is a specific adaptation to fulfill the conditions of the theory that a sensed problem must not be ignored or discarded until it has been resolved. It follows that more than one synaptic connection between two specific neurons could exist. This feature was indeed also part of the random initialization of the network as described next.
In short, if the activity of the incoming synapse is higher than that of the receiver neuron, then strengthen the weight of the synapse. If the receiver neuron activity is too high, then reduce all synaptic weights. If the activity in both the incoming synapse and the receiver is high, then disconnect that synapse from the receiver and reinsert it on some eligible receiver neuron.
Figure 5: Example implementation of the learning model based on the present conjectures. (A) The inverted pendulum implemented in the MuJoCo simulation environment Todorov et al. [2012], with the 8 sensors (S=8), angle and position plus their respective derivatives, labeled by the two alternative directions, (L)eft and (R)ight. The figure also indicates the two output actions, F+ and F-, driven by the two output cells (M=2). (B) The simple cell system configuration, where the eight sensor neurons (yellow) project to the two output neurons (red) via weighted synapses ($W_{1}$). The action selection (F- or F+) were calculated from the activity of the two output neurons as indicated. (C) The larger cell system configuration with an intermediate layer (blue), which also featured recurrent synaptic connections between all of the neurons within the layer. (D) Example data from one complete learning process using the network in B. The top panel shows the running averages of the two motor neuron activities (30 second boxcar window). Middle two panels show the running averages for the sensor activities (colored lines), and the average total sensor activity (black line), split per direction. The bottom panel shows the running average of the synaptic weights ($\pm$3 standard deviations of the synaptic change per timestep; 5 second boxcar window), as well as the running average of the topological stability of the network (i.e. the inverse of the rate of synapse movements between the target neurons; 10 second boxcar window). (E) Running grand averages of the topological stability and the total sensor activity, respectively, for N=10 simulations using the network configuration in B. (F) Similar display as in (E) but for the larger network configuration (shown in C; N=5 simulations). The topological stability was illustrated separately for the input to the I-layer and the recurrent connections within the I-layer. Both panels E and F are normalized with respect to the time until reaching a stable solution.
1.7.2 The learning process for the simpler cell system configuration
In the first example network (Figure 5B), we provided the sensor signals as weighted synaptic inputs directly to the two motor neurons. Each sensory neuron projected 25 synapses to these two eligible recipient neurons with initial randomization. Since there were only two motor neurons but eight sensory neurons there was a high degree of synaptic overlap from the beginning. The physics simulation of the inverse pendulum was started and the connectivity of the network was adapted according to the sensory minimization principle: if a synapse ($i$) is highly active ($a$) in the receiving neuron, then update the weight ($w_{i}$) of that sensor synapse. This will lead to a change in the output of the neuron ($A$) (Equation 1), which aims to decrease its own sensor input signals over time (Equation 2). Furthermore, by having two set points (both set to zero) we enabled this implementation to demonstrate the principle of sensory prioritization — the learning rule prioritizes adaptations of the most active sensor signals within each neuron (Equation 2). Adaptations consisted in weight increases for highly active sensors, or in weight decreases if the output activity of the neuron was too high (Equation 2), whereas if both the input activity and the output activity was high then the synapse was removed from that neuron and randomly assigned to another eligible neuron (including the same cell; see Appendix 1). The latter was hence an adaptation to the case where a solution could not be found with the current anatomical synapse connectivity — then the synapses with the highest activity was instead randomly re-routed. Once a sensor signal was brought down, other sensor signals would be of higher magnitude and then the neuron focused its learning to bring down also those sensor signals. This could lead to that other sensor signals rose in magnitude as a consequence of the new adaptations, resulting in a natural staging of how the solution evolved. Evidence of such phased problem solving of specific sensor inputs can be seen in the plot of the sensor signals (in Figure 5D). The next plot (Figure 5E) illustrates that the sum of all sensory inputs gradually went down while the network topology stabilized in parallel, until the solution was found and no further changes occurred. Across different random initializations, in some cases the later stages of adaptation could fail, which could lead to previously solved sub-problems reoccurring, and then the solution discovery process will move back and forth until a stable solution can be achieved.
1.7.3 The learning process for the larger cell system configuration
Using the same learning principle, we also tested a system configured with another layer of neurons between the sensors and the actuators, where all the neurons within this intermediary layer were potentially recurrently connected to each other (Figure 5C; like the sensor neurons, each intermediary neuron issued 25 synapses that were initially randomly distributed to motorneurons and other intermediary neurons). The sensory input layer only provided input to the intermediate layer (i.e. this cell population would be the equivalent of spinal interneurons Jankowska [1992], Kohler et al. [2022]), and all the cells in the intermediate layer innervated the two actuator cells. This system adapted in a similar manner as the simpler configuration, i.e. by achieving a gradually decreasing sensory input in parallel with a gradually stabilizing synaptic topology, but the solution was more stepwise led by stabilization of the recurrent connections within the intermediary population (Figure 5F). In this configuration we hence illustrated the principle that cells will propagate their unresolved problems to other cells (Conjectures 1.4-1.8), which may thus indirectly help to reduce the sensory input to that cell by utilizing downstream cells as remote enactors to resolve the problem sensed by that cell.
For both of these cell systems (Figure 5B,C), the synaptic weights evolved until a stability in the network topology was reached (Supplementary Figures 1-4). It is noteworthy that even with this bare bones model, different cell system configurations were able to solve the same problem.
1.7.4 Generalization of the learning model
The simplicity of the problem context, the inverted pendulum, provides a limited evaluation of the breadth of the applicability of the learning model. However, based on the same foundational conjectures, adaptations of this learning model can be implemented to also solve very different problems related to systems with more biological-like anatomical structures, e.g. a quadruped and a humanoid robot (Figure 6). In these cases we employed non-disclosed proprietary implementations. We illustrate the macroscopic learning processes achieved for the quadruped (real-world learning) and the biped humanoid (simulation in MuJoCo Todorov et al. [2012]) in video 1 and video 2 respectively.
Figure 6: Different environments used as contextual realization for more advanced implementations of the learning model, showcasing the potential of the current conjectures. (A) Inverted pendulum, same environment as above, with 2 motor outputs and 8 sensors. (B) Unitree Go2 quadruped robot, illustrated learning process in real world in video 1 for 12 motor outputs. (C) Unitree G1 bipedal robot, illustrated learning process for 24 motor outputs in simulation Todorov et al. [2012] in video 2.
@@ 1.8 Relation to Learning Frameworks and Plasticity
It is our hope that the framework we provide here can provide new insights into leading neuroscience theories and experimental observations on biological learning, for example why Hebbian learning and spike timing dependent plasticity are inferrable in experiments, while networks over time still arrive at highly stable homeostatic states. It should be noted that observations that unicellular organisms learn dates back more than 100 years (reviewed by Gershman et. al. Gershman et al. [2021]) and the idea that learning in the brain can be seen as a result of the need to keep the individual cells in homeostatic balance was proposed before Levin [2019]. We build on and extend those ideas towards more specific mechanistic explanations through formalizing what is being sensed, how it is being sensed and prioritized, and the passing of locally unresolved problems as internally created sensor signals to more suitable subspaces.
@@
In Artificial Intelligence, input is typically in the form of structured data, to be transformed with the intent of extracting more valuable data. By supplying large amounts of such input to these networks and utilizing training to translate that data into valuable information for a target task, the field of deep learning builds intelligent systems. We want to applaud the impressive results in that field. However, here we show through the presented conjectures that biological organisms very likely do not share the same process and most probably cannot intrinsically value raw data in the form that deep learning makes use of it. Instead, biological organisms sense only solvable, value-estimable problems that need to be surmounted to ensure survival, and this understanding of when there is a need to adapt is what allows those organisms to learn in the first place.